OpenStudy (anonymous):

Someone PLESE HELP Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

A:y = -1 over 3x - 6 B:y = -1 over 3x - 14 C:y = 3x - 6 D:y = 3x - 14

OpenStudy (anonymous):

@SolomonZelman @blurbendy @Abhisar @Tazmaniadevil

OpenStudy (anonymous):

@SithsAndGiggles @gamer56 @AquaStone316

OpenStudy (anonymous):

Someone PLEASE HELP !! I will give a MEDAL to the best answer :( <3

OpenStudy (anonymous):

I don't know, I'm pretty bad at Math. Sorry.

OpenStudy (anonymous):

Thats ok , Thanks anyways

OpenStudy (tkhunny):

Do you know the Point-Slope form of a line? You should. You WILL be required to know it on your exam.